# Stochastic gravitational wave background anisotropies from inflation with non-Bunch-Davies states

### v2

@@ 4.1 General expression for the anisotropies
It has been shown in Refs. [102, 103, 58, 104, 105, 21, 22, 23, 24, 26, 27, 28, 29, 30, 32] that the primordial non-Gaussianities result in spatial modulations of the primordial scalar and tensor power spectra. In particular, the STT bispectrum induces the directional dependence in the tensor power spectrum: [27, 29, 32]
$\mathcal{P}_{h}^{(s)}(k,\hat{n})=\bar{\mathcal{P}}^{(s)}_{h}(k)\biggl{[}1+\sum_{s^{\prime}}\int\frac{{\rm d}^{3}q}{(2\pi)^{3}}e^{id{\bf q}\cdot\hat{n}}F^{ss^{\prime}}_{\rm NL}({\bf k},{\bf q})\zeta({\bf q})\biggr{]},$ (4.1)
where $\hat{n}:=-{\bf k}/k$ and $d:=\eta_{\rm now}-\eta_{\rm ent}$ with $\eta_{\rm now}$ and $\eta_{\rm ent}$ being the conformal time at present and at the horizon reentry of inflationary GWs for each $k$ mode, respectively. Since we consider the nHz or deci-Hz inflationary GWs that re-entered the horizon long before the present time, that is, $\eta_{\rm now}\gg\eta_{\rm ent}$, we assume $d\simeq\eta_{\rm now}$. Hereinafter, a bar means the homogeneous component. In the above expression, $F^{ss^{\prime}}_{\rm NL}({\bf k},{\bf q})$ can be characterized by the cross bispectrum (3.7) as
$F^{ss^{\prime}}_{\rm NL}({\bf k},{\bf q}):=\mathcal{B}^{ss^{\prime}}_{shh}({\bf k}-{\bf q}/2,-{\bf k}-{\bf q}/2,{\bf q})|_{q\to 0}\biggl{(}\frac{\pi^{2}}{k^{3}}\mathcal{P}^{(s)}_{h}(k)\biggr{)}^{-1}\biggl{(}\frac{2\pi^{2}}{q^{3}}\mathcal{P}_{\zeta}(q)\biggr{)}^{-1},$ (4.2)
@@
${\bf k}_{1}={\bf q},~{}{\bf k}_{2}={\bf k}-{\bf q}/2,~{}{\bf k}_{3}=-{\bf k}-{\bf q}/2.$ (4.3)
Note that $k$ and $q$ in this section corresponds to $k_{L}(=k_{1})$ and $k_{S}(\simeq k_{2}\simeq k_{3})$ used in the above sections, respectively. As for the SGWB, the density parameter per logarithmic wavenumber:
$\Omega_{\rm GW}:=\frac{1}{\rho_{\rm cr}}\frac{{\rm d}\rho_{\rm GW}}{{\rm d}\ln k},$ (4.4)
has been widely used, where $\rho_{\rm cr}$ is the critical density at present, and for the inflationary GWs it is proportional to the primordial tensor power spectrum. Thus, the anisotropies of the SGWB induced from the primordial STT bispectrum can be characterized as [27, 29, 32]
$\Omega_{\rm GW}(k,\hat{n})=\bar{\Omega}_{\rm GW}(k)\biggl{[}1+\delta_{\rm GW}(k,\hat{n})\biggr{]},$ (4.5)
@@
$\delta_{\rm GW}(k,\hat{n}):=\int\frac{{\rm d}^{3}q}{(2\pi)^{3}}e^{id{\bf q}\cdot\hat{n}}\sum_{ss^{\prime}}\tilde{F}^{ss^{\prime}}_{\rm NL}({\bf k},{\bf q})\zeta({\bf q}),$ (4.6)
@@
$\tilde{F}_{\rm NL}^{ss^{\prime}}({\bf k},{\bf q}):=\frac{\bar{\mathcal{P}}^{(s)}_{h}F^{ss^{\prime}}_{\rm NL}({\bf k},{\bf q})}{\sum_{s}\bar{\mathcal{P}}^{(s)}_{h}}.$ (4.7)
@@
$\tilde{F}^{s_{2}s_{3}}_{\rm NL}({\bf k},{\bf q})=Y_{00}(\hat{n}\cdot\hat{q})f^{s_{2}s_{3}}_{\rm NL,0}+Y_{20}(\hat{n}\cdot\hat{q})f^{s_{2}s_{3}}_{\rm NL,2}+Y_{40}(\hat{n}\cdot\hat{q})f^{s_{2}s_{3}}_{\rm NL,4}+\cdots=\sum_{L={\rm even}}f^{s_{2}s_{3}}_{{\rm NL},L}Y_{L0}(\hat{n}\cdot\hat{q}),$ (4.9)
@@
×$\quad\times\int\frac{{\rm d}^{3}q}{(2\pi)^{3}}Y_{L^{\prime}M^{\prime}}(\hat{q})Y_{LM}(\hat{q})j_{L^{\prime}}(qd)\sum_{s_{2},s_{3}}f^{s_{2}s_{3}}_{{\rm NL},L}(k,q)\zeta_{\bf q},$ (4.10)
with $\begin{pmatrix}l_{1}&l_{2}&l_{3}\\ L_{1}&L_{2}&L_{3}\end{pmatrix}$ and $j_{L}(x)$ being the Wigner-$3j$ symbol and spherical Bessel function respectively.
@@ 4.2 Angular correlation for the SGWB anisotropies
$C^{\rm GW}_{\ell}(k)=\frac{1}{2\ell+1}\sum_{m}\langle\delta^{\rm GW}_{\ell m}(k)\delta^{\rm GW*}_{\ell m}(k)\rangle=\sum_{L={\rm even}}C^{\rm GW}_{\ell,L}(k),$ (4.11)
@@
×$\quad\times\int\frac{{\rm d}q}{q}j_{J}(qd)j_{J^{\prime}}(qd)\mathcal{P}_{\zeta}(q)\sum_{s_{2},s_{3},s^{\prime}_{2},s^{\prime}_{3}}f^{s_{2}s_{3}}_{{\rm NL},L}(k,q)f^{s^{\prime}_{2}s^{\prime}_{3}}_{{\rm NL},L}(k,q).$ (4.12)
@@
$a^{X}_{\ell m}=4\pi i^{\ell}\int\frac{{\rm d}^{3}p}{(2\pi)^{3}}\mathcal{T}^{X}_{\ell}(p)Y^{*}_{\ell m}(\hat{p})\zeta({\bf p}),$ (4.14)
where the $X$ stands for $T$ (temperature fluctuations) or $E$ (E-mode polarization). [^3] As has been studied in Ref. [32], the cross-correlation between the B-mode polarization and the SGWB anisotropy sourced by the long-wavelength tensor mode exists in parity-violating gravitational theories. In the same way as the angular auto-correlation function of the SGWB anisotropies, the cross-correlation function of the CMB and SGWB anisotropies can be expressed as
$C^{\rm GW-X}_{\ell}(k)=\frac{1}{2\ell+1}\sum_{m}\langle\delta^{\rm GW}_{\ell m}(k)a^{X*}_{\ell m}\rangle=\sum_{L={\rm even}}C^{\rm GW-X}_{\ell,L}(k),$ (4.15)
@@
$C^{\rm GW-X}_{\ell,L}(k)=4\pi\sqrt{\frac{2L+1}{4\pi}}\sum_{J}i^{J-\ell}(2J+1)\biggl{[}\begin{pmatrix}\ell&J&L\\ 0&0&0\end{pmatrix}\biggr{]}^{2}\int\frac{{\rm d}q}{q}j_{J}(qd)\mathcal{T}^{X}_{\ell}(q)\mathcal{P}_{\zeta}(q)\sum_{s_{2},s_{3}}f^{s_{2}s_{3}}_{{\rm NL},L}(k,q),$ (4.16)
@@ 4.3 SGWB anisotropies from inflation with the non-Bunch-Davies states
Then, let us evaluate the angular correlation function sourced from the primordial scalar-tensor-tensor bispectrum with the non-Bunch-Davies states. In the present setup, one can take the squeezed limit to the STT bispectrum (3.10) straightforwardly as
@@
×$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\times(\hat{n}\cdot\hat{q})\times c_{h}k\eta_{0}\times c_{s}q\eta_{0}+O((q/k)^{2}).$ (4.17)
Note that here we assume $s_{2}=s_{3}=s$. This is because as can be seen in the expressions for $\mathcal{V}_{i}$ at the squeezed limit studied in Appendix C, $\mathcal{V}_{i}$ are suppressed at most in the order of $q/k$ for the case with $s_{2}=-s_{3}$. Thus, in the leading order, not only the power spectrum but also the bispectrum depend on the choice of the helicity only through the helicity dependence of the Bogoliubov coefficients. In the following subsections, we will investigate the correlation functions of the SGWB anisotropies for the two cases of the Bogoliubov coefficients which are introduced in section 2.2.
@@ 4.3.1 Peaked spectrum
The fourth line in the squeezed bispectrum  (4.17), which is proportional to $\hat{n}\cdot\hat{q}$, has
$(\beta^{(s)*}_{k_{2}}\alpha^{(s)*}_{k_{3}}-\alpha^{(s)*}_{k_{2}}\beta^{(s)*}_{k_{3}}).$ (4.18)
By substituting the assumptions for the Bogoliubov coefficients given by Eqs. (2.31) and (2.32), one can find that it is reduced to ${\rm Re}[\beta^{(s)}_{k_{2}}]-{\rm Re}[\beta^{(s)}_{k_{3}}]$ and at the leading order in the squeezed limit this should be the order of $(\hat{n}\cdot\hat{q})q$. Thus, in the squeezed limit the fourth line of Eq. (4.17) is found to be proportional to $(\hat{n}\cdot\hat{q})^{2}\times(c_{s}q\eta_{0})^{2}$. Since we focus on the phase after the long-wavelength mode $q$ crosses the sound horizon, i.e., the phase characterized by $|c_{s}q\eta_{0}|\ll 1$, there is no enhancement for the term proportional to $(\hat{n}\cdot\hat{q})^{2}$ in the bispectrum. Accordingly, it is enough to consider the terms proportional to $(\hat{n}\cdot\hat{q})^{0}$.
From Eqs. (2.31) and (2.32), up to the leading order in $q/k$, we can also find
${\rm Re}$ $\left[\frac{(\alpha^{(s)}_{k_{2}}-\beta^{(s)}_{k_{2}})(\alpha^{(s)}_{k_{3}}-\beta^{(s)}_{k_{3}})(\alpha^{(s)*}_{k_{2}}\alpha^{(s)*}_{k_{3}}-\beta^{(s)*}_{k_{2}}\beta^{(s)*}_{k_{3}})}{|\alpha^{(s)}_{k}-\beta^{(s)}_{k}|^{2}}\right]\approx 1,$ (4.19)
@@
$\qquad\qquad\qquad\approx\frac{2A^{2}}{2A^{2}-2Ae^{\ln^{2}(k/k_{*})/\Delta^{2}}+e^{2\ln^{2}(k/k_{*})/\Delta^{2}}}=:\mathcal{A}(k,k_{*}).$ (4.20)
For $k\approx k_{*}$ and $A\gg 1$, $\mathcal{A}(k,k_{*})\approx 1$. As a result, the non-linearity parameter in the squeezed limit of the scalar-tensor-tensor bispectrum $F_{\rm NL}^{s}$ for the peaked tensor power spectrum can be rewritten as
$\tilde{F}_{\rm NL}^{ss}:=\frac{\mathcal{B}^{ss}_{shh}|_{q\to 0}}{P_{s}(q)P_{h}^{(s)}(k)}$ $=\left(\frac{4b_{1}}{\mathcal{G}_{T}}+\frac{12b_{2}}{\mathcal{F}_{T}}\right)-\left[8\left(\frac{b_{1}}{\mathcal{G}_{T}}+\frac{b_{2}}{\mathcal{F}_{T}}\right)c_{h}k\eta_{0}-\frac{8b_{2}}{\mathcal{F}_{T}}\frac{1}{c_{h}k\eta_{0}}\right]\mathcal{A}(k,k_{*}),$ (4.21)
@@
$f^{ss}_{{\rm NL},0}(k)=\frac{\tilde{F}^{ss}_{\rm NL}}{Y_{00}}=2\sqrt{\pi}\biggl{\{}\left(\frac{4b_{1}}{\mathcal{G}_{T}}+\frac{12b_{2}}{\mathcal{F}_{T}}\right)-\left[8\left(\frac{b_{1}}{\mathcal{G}_{T}}+\frac{b_{2}}{\mathcal{F}_{T}}\right)c_{h}k\eta_{0}-\frac{8b_{2}}{\mathcal{F}_{T}}\frac{1}{c_{h}k\eta_{0}}\right]\mathcal{A}(k,k_{*})\biggr{\}}.$ (4.22)
@@
We then compute the correlation functions of the SGWB anisotropies. From Eq. (4.12), we obtain
$C^{\rm GW}_{\ell}(k)=\frac{4\mathcal{P}_{\zeta}f^{ss}_{{\rm NL},0}(k)^{2}}{2\ell(\ell+1)},$ (4.23)
@@
$\int^{\infty}_{0}\frac{{\rm d}x}{x}j_{\ell}(x)j_{\ell}(x)=\frac{1}{2\ell(\ell+1)}.$ (4.24)
@@
$\mathcal{T}^{T}_{\ell}(k)\simeq\frac{1}{5}j_{\ell}(k(\eta_{\rm now}-\eta_{\rm dec})),$ (4.25)
@@
$C^{\rm GW-T}_{\ell}(k)\simeq\frac{4\sqrt{\pi}}{5}\mathcal{P}_{\zeta}f^{ss}_{{\rm NL},0}(k)\int\frac{{\rm d}q}{q}j_{\ell}(qd)j_{\ell}(q(\eta_{\rm now}-\eta_{\rm dec})),$ (4.26)
@@
$=\frac{1}{4}c^{\ell}\sqrt{\pi}\Gamma(\ell)\frac{{}_{2}F_{1}\biggl{(}-\frac{1}{2},\ell,\frac{3}{2}+\ell;c^{2}\biggr{)}}{\Gamma(\frac{3}{2}+\ell)},$ (4.27)
@@
$c:=\frac{\eta_{\rm now}-\eta_{\rm dec}}{d}\simeq\frac{\eta_{\rm now}-\eta_{\rm dec}}{\eta_{\rm now}}.$ (4.28)
Note that by taking $c\to 1$ to Eq. (4.27), we obtain the same result with Eq. (4.24). The resultant form of $C_{\ell}^{{\rm GW}-T}$ reads
$C_{\ell}^{\rm GW-T}(k)\simeq\frac{\pi}{5}\mathcal{P}_{\zeta}f^{ss}_{{\rm NL},0}(k)\times c^{\ell}\frac{\Gamma(\ell)}{\Gamma(\ell+\frac{3}{2})}{}_{2}F_{1}\biggl{(}-\frac{1}{2},\ell,\frac{3}{2}+\ell;c^{2}\biggr{)}.$ (4.29)
@@
$\frac{b_{1}}{\mathcal{G}_{T}}$ $\leq\mathcal{O}(10^{5})\left(\frac{|\zeta(\eta_{0})|}{10^{-5}}\right)^{-1},$ (4.30)
$\frac{b_{2}}{\mathcal{F}_{T}}$ $\leq\mathcal{O}(10^{5})\left(\frac{|\zeta(\eta_{0})|}{10^{-5}}\right)^{-1}.$ (4.31)
@@
In light of that the curvature perturbation is on the superhorizon scales at $\eta=\eta_{0}$, we can use the current constraint on the amplitude of $\zeta$ as $|\zeta(\eta_{0})|\sim 10^{-5}$ [106]. Another constraint is obtained from an argument on backreaction from the excited tensor modes. Once the modes get excited from the Bunch-Davies state, the excited modes cause backreaction to the inflationary background [107]. In the Horndeski theory, the excited tensor mode has the following energy density [67]:
$\rho_{\rm excited}=\frac{c_{h}}{a^{4}(\eta)}\int|\beta^{(s)}_{k}|^{2}k^{3}{\rm d}k.$ (4.32)
@@
In order for the backreaction from the excited modes not to spoil the background dynamics of the slow-roll inflation, we compare this to the homogeneous part of the Friedmann equation $\mathcal{E}_{b}$. In the standard slow-roll inflation, the homogeneous part corresponds to the energy density of inflaton as $\mathcal{E}_{b}\sim M_{\rm pl}^{2}H_{\rm inf}^{2}$. In the general inflation model considered here, we similarly assume $\mathcal{E}_{b}\sim M_{\rm pl}^{2}H_{\rm inf}^{2}$ for simplicity. Then, the backreaction constraint, i.e., $\rho_{\rm excited}\lesssim\mathcal{E}_{b}$, reads
$\frac{1}{a(\eta_{0})^{4}}A^{2}e^{2\Delta^{2}}k_{*}^{4}\Delta\lesssim M_{\rm pl}^{2}H_{\rm inf}^{2}.$ (4.33)
If we assume $A=-3000$, $\Delta=0.45$ and $H_{\rm inf}=10^{13}~{}{\rm GeV}$ to explain the recent PTA results, the above reads
$-k_{*}\eta_{0}\lesssim 10.$ (4.34)
@@
From Eqs. (4.30), (4.31), and (4.34), we can estimate the theoretical upper bound on $f_{\rm NL}$ as
$f^{ss}_{{\rm NL},0}(k_{*})\lesssim 10^{7}.$ (4.35)

### v3

@@ 4.1 General expression for the anisotropies
It has been shown in Refs. Jeong:2012df ; Dai:2013kra ; Bahrami:2013isa ; Dimastrogiovanni:2014ina ; Dimastrogiovanni:2015pla ; Ricciardone:2017kre ; Dimastrogiovanni:2018uqy ; Dimastrogiovanni:2019bfl ; Bartolo:2019oiq ; Bartolo:2019yeu ; Adshead:2020bji ; Malhotra:2020ket ; Dimastrogiovanni:2021mfs ; Dimastrogiovanni:2022afr ; Orlando:2022rih that the primordial non-Gaussianities result in spatial modulations of the primordial scalar and tensor power spectra. In particular, the STT bispectrum induces the directional dependence in the tensor power spectrum: Adshead:2020bji ; Dimastrogiovanni:2021mfs ; Orlando:2022rih
$\mathcal{P}_{h}^{(s)}(k,\hat{n})=\bar{\mathcal{P}}^{(s)}_{h}(k)\biggl[1+\sum_{s^{\prime}}\int\frac{{\rm d}^{3}q}{(2\pi)^{3}}e^{id{\bf q}\cdot\hat{n}}F^{ss^{\prime}}_{\rm NL}({\bf k},{\bf q})\zeta({\bf q})\biggr],$ (83)
where $\hat{n}:=-{\bf k}/k$ and $d:=\eta_{\rm now}-\eta_{\rm ent}$ with $\eta_{\rm now}$ and $\eta_{\rm ent}$ being the conformal time at present and at the horizon reentry of inflationary GWs for each $k$ mode, respectively. Since we consider the nHz or deci-Hz inflationary GWs that reentered the horizon long before the present time, i.e. $\eta_{\rm now}\gg\eta_{\rm ent}$, we assume $d\simeq\eta_{\rm now}$. Hereinafter, a bar means the homogeneous component. In the above expression, $F^{ss^{\prime}}_{\rm NL}({\bf k},{\bf q})$ can be characterized by the cross bispectrum (46) as
$F^{ss^{\prime}}_{\rm NL}({\bf k},{\bf q}):=\mathcal{B}^{ss^{\prime}}_{shh}({\bf k}-{\bf q}/2,-{\bf k}-{\bf q}/2,{\bf q})|_{q\to 0}\biggl(\frac{\pi^{2}}{k^{3}}\mathcal{P}^{(s)}_{h}(k)\biggr)^{-1}\biggl(\frac{2\pi^{2}}{q^{3}}\mathcal{P}_{\zeta}(q)\biggr)^{-1},$ (84)
@@
${\bf k}_{1}={\bf q},~{\bf k}_{2}={\bf k}-{\bf q}/2,~{\bf k}_{3}=-{\bf k}-{\bf q}/2.$ (85)
Note that $k$ and $q$ in this section corresponds to $k_{L}(=k_{1})$ and $k_{S}(\simeq k_{2}\simeq k_{3})$ used in the above sections, respectively. As for the SGWB, the density parameter per logarithmic wavenumber,
$\Omega_{\rm GW}:=\frac{1}{\rho_{\rm cr}}\frac{{\rm d}\rho_{\rm GW}}{{\rm d}\ln k},$ (86)
has been widely used, where $\rho_{\rm cr}$ is the critical density at present, and for the inflationary GWs it is proportional to the primordial tensor power spectrum. Thus, the anisotropies of the SGWB induced from the primordial STT bispectrum can be characterized as Adshead:2020bji ; Dimastrogiovanni:2021mfs ; Orlando:2022rih :
[…]
@@
$\delta_{\rm GW}(k,\hat{n}):=\int\frac{{\rm d}^{3}q}{(2\pi)^{3}}e^{id{\bf q}\cdot\hat{n}}\sum_{ss^{\prime}}\tilde{F}^{ss^{\prime}}_{\rm NL}({\bf k},{\bf q})\zeta({\bf q}),$ (88)
@@
$\tilde{F}_{\rm NL}^{ss^{\prime}}({\bf k},{\bf q}):=\frac{\bar{\mathcal{P}}^{(s)}_{h}F^{ss^{\prime}}_{\rm NL}({\bf k},{\bf q})}{\sum_{s}\bar{\mathcal{P}}^{(s)}_{h}}$. (89)
@@
[…]
@@
$\quad\times\int\frac{{\rm d}^{3}q}{(2\pi)^{3}}Y_{L^{\prime}M^{\prime}}(\hat{q})Y_{LM}(\hat{q})j_{L^{\prime}}(qd)\sum_{s_{2},s_{3}}f^{s_{2}s_{3}}_{{\rm NL},L}(k,q)\zeta_{\bf q},$ (92)
with $\begin{pmatrix}l_{1}&l_{2}&l_{3}\\ L_{1}&L_{2}&L_{3}\end{pmatrix}$ and $j_{L}(x)$ being the Wigner-$3j$ symbol and spherical Bessel function, respectively.
@@ 4.2 Angular correlation for the SGWB anisotropies
$C^{\rm GW}_{\ell}(k)=\frac{1}{2\ell+1}\sum_{m}\langle\delta^{\rm GW}_{\ell m}(k)\delta^{\rm GW*}_{\ell m}(k)\rangle=\sum_{L={\rm even}}C^{\rm GW}_{\ell,L}(k),$ (93)
@@
$\quad\times\int\frac{{\rm d}q}{q}j_{J}(qd)j_{J^{\prime}}(qd)\mathcal{P}_{\zeta}(q)\sum_{s_{2},s_{3},s^{\prime}_{2},s^{\prime}_{3}}f^{s_{2}s_{3}}_{{\rm NL},L}(k,q)f^{s^{\prime}_{2}s^{\prime}_{3}}_{{\rm NL},L}(k,q).$ (94)
@@
$a^{X}_{\ell m}=4\pi i^{\ell}\int\frac{{\rm d}^{3}p}{(2\pi)^{3}}\mathcal{T}^{X}_{\ell}(p)Y^{*}_{\ell m}(\hat{p})\zeta({\bf p}),$ (96)
where the $X$ stands for $T$ (temperature fluctuations) or $E$ (E-mode polarization). [^3] As has been studied in Ref. Orlando:2022rih , the cross-correlation between the B-mode polarization and the SGWB anisotropy sourced by the long-wavelength tensor mode exists in parity-violating gravitational theories. In the same way as the angular auto-correlation function of the SGWB anisotropies, the cross-correlation function of the CMB and SGWB anisotropies can be expressed as
$C^{\rm GW-X}_{\ell}(k)=\frac{1}{2\ell+1}\sum_{m}\langle\delta^{\rm GW}_{\ell m}(k)a^{X*}_{\ell m}\rangle=\sum_{L={\rm even}}C^{\rm GW-X}_{\ell,L}(k),$ (97)
@@
$C^{\rm GW-X}_{\ell,L}(k)=4\pi\sqrt{\frac{2L+1}{4\pi}}\sum_{J}i^{J-\ell}(2J+1)\biggl[\begin{pmatrix}\ell&J&L\\ 0&0&0\end{pmatrix}\biggr]^{2}\int\frac{{\rm d}q}{q}j_{J}(qd)\mathcal{T}^{X}_{\ell}(q)\mathcal{P}_{\zeta}(q)\sum_{s_{2},s_{3}}f^{s_{2}s_{3}}_{{\rm NL},L}(k,q),$ (98)
@@ 4.3 SGWB anisotropies from inflation with the non-Bunch-Davies states
Then, let us evaluate the angular correlation function sourced from the primordial STT bispectrum with the non-Bunch-Davies states. In the present setup, one can take the squeezed limit to the STT bispectrum (49) straightforwardly as
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\times(\hat{n}\cdot\hat{q})\times c_{h}k\eta_{0}\times c_{s}q\eta_{0}+O((q/k)^{2}).$ (99)
Note that here we assume $s_{2}=s_{3}=s$. This is because, as can be seen in the expressions for $\mathcal{V}_{i}$ at the squeezed limit studied in Appendix C, $\mathcal{V}_{i}$ are suppressed at most in the order of $q/k$ for the case with $s_{2}=-s_{3}$. Thus, in the leading order, not only the power spectrum but also the bispectrum depend on the choice of the helicity only through the helicity dependence of the Bogoliubov coefficients. In the following subsections, we will investigate the correlation functions of the SGWB anisotropies for the two cases of the Bogoliubov coefficients which are introduced in section 2.2.
@@ 4.3.1 Peaked spectrum
The fourth line in the squeezed bispectrum  (99), which is proportional to $\hat{n}\cdot\hat{q}$, has
$(\beta^{(s)*}_{k_{2}}\alpha^{(s)*}_{k_{3}}-\alpha^{(s)*}_{k_{2}}\beta^{(s)*}_{k_{3}})$. (100)
By substituting the assumptions for the Bogoliubov coefficients given by Eqs. (31) and (32), one can find that it is reduced to ${\rm Re}[\beta^{(s)}_{k_{2}}]-{\rm Re}[\beta^{(s)}_{k_{3}}$] and at the leading order in the squeezed limit this should be on the order of $(\hat{n}\cdot\hat{q})q$. Thus, in the squeezed limit the fourth line of Eq. (99) is found to be proportional to $(\hat{n}\cdot\hat{q})^{2}\times(c_{s}q\eta_{0})^{2}$. Since we focus on the phase after the long-wavelength mode $q$ crosses the sound horizon, i.e. the phase characterized by $|c_{s}q\eta_{0}|\ll 1$, there is no enhancement for the term proportional to $(\hat{n}\cdot\hat{q})^{2}$ in the bispectrum. Accordingly, it is enough to consider the terms proportional to $(\hat{n}\cdot\hat{q})^{0}$.
From Eqs. (31) and (32), up to the leading order in $q/k$, we can also find
${\rm Re}$ $\left[\frac{(\alpha^{(s)}_{k_{2}}-\beta^{(s)}_{k_{2}})(\alpha^{(s)}_{k_{3}}-\beta^{(s)}_{k_{3}})(\alpha^{(s)*}_{k_{2}}\alpha^{(s)*}_{k_{3}}-\beta^{(s)*}_{k_{2}}\beta^{(s)*}_{k_{3}})}{|\alpha^{(s)}_{k}-\beta^{(s)}_{k}|^{2}}\right]\approx 1,$ (101)
@@
$\qquad\qquad\qquad\approx\frac{2A^{2}}{2A^{2}-2Ae^{\ln^{2}(k/k_{*})/\Delta^{2}}+e^{2\ln^{2}(k/k_{*})/\Delta^{2}}}=:\mathcal{A}(k,k_{*}).$ (102)
For $k\approx k_{*}$ and $A\gg 1$, $\mathcal{A}(k,k_{*})\approx 1$. One can also find $\mathcal{A}\ll 1$ for $A\ll 1$ which suppresses terms in the non-linearity parameter newly induced by non-Bunch-Davies effects as shown below. Therefore, $A\gg 1$ is found to be important to enhance both SGWBs and anisotropies. As a result, the non-linearity parameter in the squeezed limit of the STT bispectrum $F_{\rm NL}^{s}$ for the peaked tensor power spectrum can be rewritten as
$\tilde{F}_{\rm NL}^{ss}:=\frac{\mathcal{B}^{ss}_{shh}|_{q\to 0}}{P_{s}(q)P_{h}^{(s)}(k)}$ $=\left(\frac{4b_{1}}{\mathcal{G}_{T}}+\frac{12b_{2}}{\mathcal{F}_{T}}\right)-\left[8\left(\frac{b_{1}}{\mathcal{G}_{T}}+\frac{b_{2}}{\mathcal{F}_{T}}\right)c_{h}k\eta_{0}-\frac{8b_{2}}{\mathcal{F}_{T}}\frac{1}{c_{h}k\eta_{0}}\right]\mathcal{A}(k,k_{*}),$ (103)
@@
$f^{ss}_{{\rm NL},0}(k)=\frac{\tilde{F}^{ss}_{\rm NL}}{Y_{00}}=2\sqrt{\pi}\biggl\{\left(\frac{4b_{1}}{\mathcal{G}_{T}}+\frac{12b_{2}}{\mathcal{F}_{T}}\right)-\left[8\left(\frac{b_{1}}{\mathcal{G}_{T}}+\frac{b_{2}}{\mathcal{F}_{T}}\right)c_{h}k\eta_{0}-\frac{8b_{2}}{\mathcal{F}_{T}}\frac{1}{c_{h}k\eta_{0}}\right]\mathcal{A}(k,k_{*})\biggr\}.$ (104)
@@
We then compute the correlation functions of the SGWB anisotropies. From Eq. (94), we obtain
$C^{\rm GW}_{\ell}(k)=\frac{4\mathcal{P}_{\zeta}f^{ss}_{{\rm NL},0}(k)^{2}}{2\ell(\ell+1)},$ (105)
@@
$\int^{\infty}_{0}\frac{{\rm d}x}{x}j_{\ell}(x)j_{\ell}(x)=\frac{1}{2\ell(\ell+1)}.$ (106)
@@
$\mathcal{T}^{T}_{\ell}(k)\simeq\frac{1}{5}j_{\ell}(k(\eta_{\rm now}-\eta_{\rm dec})),$ (107)
@@
$C^{\rm GW-T}_{\ell}(k)\simeq\frac{4\sqrt{\pi}}{5}\mathcal{P}_{\zeta}f^{ss}_{{\rm NL},0}(k)\int\frac{{\rm d}q}{q}j_{\ell}(qd)j_{\ell}(q(\eta_{\rm now}-\eta_{\rm dec})),$ (108)
@@
$=\frac{1}{4}c^{\ell}\sqrt{\pi}\Gamma(\ell)\frac{{}_{2}F_{1}\biggl(-\frac{1}{2},\ell,\frac{3}{2}+\ell;c^{2}\biggr)}{\Gamma(\frac{3}{2}+\ell)},$ (109)
@@
$c:=\frac{\eta_{\rm now}-\eta_{\rm dec}}{d}\simeq\frac{\eta_{\rm now}-\eta_{\rm dec}}{\eta_{\rm now}}.$ (110)
Note that by taking $c\to 1$ to Eq. (109), we obtain the same result with Eq. (106). The resultant form of $C_{\ell}^{{\rm GW}-T}$ reads
$C_{\ell}^{\rm GW-T}(k)\simeq\frac{\pi}{5}\mathcal{P}_{\zeta}f^{ss}_{{\rm NL},0}(k)\times c^{\ell}\frac{\Gamma(\ell)}{\Gamma(\ell+\frac{3}{2})}{}_{2}F_{1}\biggl(-\frac{1}{2},\ell,\frac{3}{2}+\ell;c^{2}\biggr).$ (111)
@@
$\frac{b_{1}}{\mathcal{G}_{T}}$ $\leq\mathcal{O}(10^{5})\left(\frac{|\zeta(\eta_{0})|}{10^{-5}}\right)^{-1},$ (112)
$\frac{b_{2}}{\mathcal{F}_{T}}$ $\leq\mathcal{O}(10^{5})\left(\frac{|\zeta(\eta_{0})|}{10^{-5}}\right)^{-1}.$ (113)
@@
In light of the fact that the curvature perturbation is on the superhorizon scales at $\eta=\eta_{0}$, we can use the current constraint on the amplitude of $\zeta$ as $|\zeta(\eta_{0})|\sim 10^{-5}$ Planck:2018jri . Another constraint is obtained from an argument on backreaction from the excited tensor modes. Once the modes get excited from the Bunch-Davies state, the excited modes cause backreaction to the inflationary background Tanaka:2000jw . In the Horndeski theory, the excited tensor mode has the following energy density Akama:2020jko :
$\rho_{\rm excited}=\frac{c_{h}}{a^{4}(\eta)}\int|\beta^{(s)}_{k}|^{2}k^{3}{\rm d}k.$ (114)
@@
In order for the backreaction from the excited modes not to spoil the background dynamics of the slow-roll inflation, we compare this to the homogeneous part of the Friedmann equation $\mathcal{E}_{b}$. In the standard slow-roll inflation, the homogeneous part corresponds to the energy density of the inflaton as $\mathcal{E}_{b}\sim M_{\rm pl}^{2}H_{\rm inf}^{2}$. In the general inflation model considered here, we similarly assume $\mathcal{E}_{b}\sim M_{\rm pl}^{2}H_{\rm inf}^{2}$ for simplicity. Then, the backreaction constraint, i.e. $\rho_{\rm excited}\lesssim\mathcal{E}_{b}$, reads
$\frac{1}{a(\eta_{0})^{4}}A^{2}e^{2\Delta^{2}}k_{*}^{4}\Delta\lesssim M_{\rm pl}^{2}H_{\rm inf}^{2}.$ (115)
If we assume A=3000A=-3000, $\Delta=0.45$, and $H_{\rm inf}=10^{13}~{\rm GeV}$ to explain the recent PTA results, the above reads
$-k_{*}\eta_{0}\lesssim 10,$ (116)
@@
which is consistent with Eq. (45). From Eqs. (112), (113), and (116), we can estimate the theoretical upper bound on $f_{\rm NL}$ as
$f^{ss}_{{\rm NL},0}(k)\lesssim 10^{7}.$ (117)